OpenStudy (yacoub1993):

As 390 g of hot milk cools in a mug, it transfers 30,000 J of heat to the environment. What is the temperature change of the milk? The specific heat of milk is 3.9 J/g•°C. Show your work.

OpenStudy (yacoub1993):

@ybarrap i need help here please

OpenStudy (yacoub1993):

i got it bro i tried my best and found it out thnxs bro @ybarrap

OpenStudy (yacoub1993):

delta T = q/mc delta T is change in temperature, q is the quantity of heat transferred, m=the mass of the substance, and c is the specific heat capacity of the substance. m = 390, q = -30000, c = 3.9 delta T = -30000/(390)(3.9) delta T = -19.7 So the milk's temperature decrease by 19.7 degrees Celcius this is what i got thnxs alot

OpenStudy (ybarrap):

I am not familiar with this type of question. However, given the units for specific heat, I see some of your givens there: Joules and Mass. I also see something you want, "C." Based on what you wrote, just make sure the unit comes out to Celcius. But I would confirm with your book that this you don't need to worry about some law or other factor.